(D^2-3D+2)y= e^x/1+e^x

1
Expert's answer
2022-01-17T16:12:03-0500


Characteristic equation

"{k^2} - 3k + 2 = 0"

"(k - 1)(k - 2) = 0"

"{k_1} = 1;\,{k_2} = 2"

We obtain the general solution of the homogeneous equation

"{y_0} = {C_1}{e^x} + {C_2}{e^{2x}}"

Let "{C_1} = f(x),\,\,{C_2} = g(x)".

Then

"y = f{e^x} + g{e^{2x}} \Rightarrow y' = f'{e^x} + f{e^x} + g'{e^{2x}} + 2g{e^{2x}}"

Let

"f'{e^x} + g'{e^{2x}} = 0\,\,\,(*)"

Then

"y' = f{e^x} + 2g{e^{2x}} \Rightarrow y'' = f'{e^x} + f{e^x} + 2g'{e^{2x}} + 4g{e^{2x}}"

Substitute in the original equation

"f'{e^x} + f{e^x} + 2g'{e^{2x}} + 4g{e^{2x}} - 3\left( {f{e^x} + 2g{e^{2x}}} \right) + 2\left( {f{e^x} + g{e^{2x}}} \right) = \frac{{{e^x}}}{{1 + {e^x}}}"

"f'{e^x} + 2g'{e^{2x}} = \frac{{{e^x}}}{{1 + {e^x}}}"

Together with (*) we have a system of equations

"\left\{ \begin{array}{l}
f'{e^x} + 2g'{e^{2x}} = \frac{{{e^x}}}{{1 + {e^x}}}\\
f'{e^x} + g'{e^{2x}} = 0
\end{array} \right."

from 2-nd eauation:

"f' = - g'{e^x}"

Substitute in the 1-st equation

"- g'{e^{2x}} + 2g'{e^{2x}} = \frac{{{e^x}}}{{1 + {e^x}}}"

"g' = \frac{{{e^{ - x}}}}{{1 + {e^x}}}"

"g = \int {\frac{{{e^{ - x}}}}{{1 + {e^x}}}dx = - x - {e^{ - x}}} + \ln ({e^x} + 1) + {C_1}"

"f' = - g'{e^x} = - \frac{1}{{1 + {e^x}}} \Rightarrow f = - \int {\frac{1}{{1 + {e^x}}}} dx = \ln ({e^x} + 1) - x + {C_2}"

Then

"y = f{e^x} + g{e^{2x}} = \left( {\ln ({e^x} + 1) - x + {C_2}} \right){e^x} + \left( { - x - {e^{ - x}} + \ln ({e^x} + 1) + {C_1}} \right){e^{2x}}"

Answer: "y = \left( {\ln ({e^x} + 1) - x + {C_2}} \right){e^x} + \left( { - x - {e^{ - x}} + \ln ({e^x} + 1) + {C_1}} \right){e^{2x}}"